# ABD: Default–Exception Abduction in Finite First-Order Worlds

### v1

@@ 6.5 Complexity and Generalization
Gemini-3 93 +1.3 60 +1.0 52 +0.6
Grok4.1f 99 +1.5 62 +0.6 86 +0.2
DSR 97 +1.5 51 +0.8 67 +0.2
Kimi-K2t 100 +1.5 86 +0.4 88 +1.8
@@
Gemini-3 80 +1.1 32 +0.5 27 +0.6
Grok4.1f 93 +1.2 36 +0.2 0
DSR 89 +1.1 73 +1.2
Kimi-K2t 95 +1.2 31 +0.9
@@
FULL 140 92 +1.52 140 49 +0.60
PARTIAL 135 84 +1.17 135 26 +0.48
SKEPTICAL 121 79 -0.00 121 32 +0.33
Overall 396 85 +0.94 396 36 +0.50
@@
Gemini-3 11.5% 0.6 24.1 53
Grok4.1f 6.5% 0.7 20.5 35
DSR 4.4% 0.8 16.8 25
Kimi-K2t 3.4% 0.7 16.1 14
@@
Table 11 shows that some models beat the planted gold rule on a nontrivial fraction of instances, confirming that gold rules are generator anchors (Section 4.4). This diagnostic identifies instances where the planted rule may be suboptimal and feeds back into generator refinement.
@@ 6.6 Discussion
Models differ qualitatively in their performance profiles. Some models (e.g., Gemini 3.1 Pro, DeepSeek Reasoner and Grok 4.1 Fast) excel in validity, while others (e.g., Gemini 3 Pro, Grok 4) achieve lower gaps among valid predictions. GPT-5.2 achieves strong overall performance on training worlds but relies on large formulas that often represent case splits. Taken together, the results support three conclusions.
@@
Parsimony remains challenging: the best systems still incur about one extra abnormality per world beyond a solver lower bound, and this gap roughly doubles on holdouts in ABD-Full/Partial.
@@
Training metrics alone are misleading because training sets are adversarially filtered. Holdouts reveal that the answer differs by scenario: parsimony inflation dominates in ABD-Full/Partial, while validity brittleness dominates in ABD-Skeptical, suggesting that robust semantics regularizes cost but makes validity harder to preserve.
@@
Binary validity saturates; cost gaps reveal separation; but gap can be driven down by hypothesis complexity. Improved abductive reasoning involves improvement in all three metrics: validity, gap cost, and formula size.
@@ 7 Conclusion
We introduced ABD, a solver-checkable benchmark suite for default–exception abduction over finite first-order worlds. In ABD, models must synthesize a first-order abnormality definition $\alpha(x)$ such that $Ab(x)\leftrightarrow\alpha(x)$ repairs a fixed default theory across multiple worlds, while keeping exceptions sparse. We formalized three observation regimes—ABD-Full (closed-world), ABD-Partial (existential completion with best-case cost), and ABD-Skeptical (universal completion with worst-case cost)—and evaluated ten LLMs under exact SMT-based verification.
Across tasks, frontier models often produce valid repairs on training worlds, but ABD remains unsaturated because parsimony does not saturate. Among strong models, the normalized gap to the solver lower bound remains roughly 1.01.01.61.6 extra abnormalities per world; the gap to the planted gold rule is roughly 0.50.51.01.0, confirming systematic over-use of exceptions. The benchmark was calibrated to be moderately challenging for frontier models; harder instances can readily be generated in the future.
@@
Finally, the complexity analyses emphasize that performance cannot be summarized by validity and unbudgeted cost. Large formulas can reduce cost on training worlds but become brittle on holdouts, consistent with case-splitting. Moderate-size formulas often provide the best tradeoff. This motivates reporting ABD performance as a multi-objective profile over validity, parsimony, and syntactic budget.
@@ G.1 Full Per-Theory Results
Opus-4.6 Grok4 GPT-5.2 Gemini-3.1 Gemini-3 Grok4.1f DSR Kimi-K2t Hermes4 GPT-4o
T1 1.3/3.2 (4) 1.1/3.2 (3) 1.3/3.7 (3) 1.5/3.6 (4) 1.1/1.9 (3) 1.4/3.2 (5) 1.5/3.3 (5) 1.4/3.1 (4)
T2 1.4/3.3 (56) 0.9/1.6 (35) 1.1/3.2 (37) 1.4/3.2 (54) 1.2/2.8 (38) 1.4/3.0 (49) 1.6/3.6 (48) 1.4/3.5 (33) 2.0/3.7 (6)
T3 1.4/2.8 (37) 1.1/2.5 (29) 0.9/2.6 (19) 1.3/2.6 (39) 1.4/2.6 (21) 1.6/2.9 (39) 1.6/2.9 (36) 1.6/3.0 (24) 1.4/3.0 (4)
T4 1.3/2.5 (46) 0.9/1.4 (27) 0.9/2.5 (28) 1.4/2.2 (44) 1.4/2.2 (25) 1.4/2.4 (49) 1.5/2.7 (45) 1.7/2.6 (24) 5.3/— (0)
T5 0.8/1.8 (27) 0.7/1.8 (22) 0.6/2.2 (18) 0.9/2.0 (27) 0.9/2.0 (25) 0.9/2.1 (24) 1.0/2.1 (26) 1.3/2.4 (24) 2.9/4.3 (6) 5.8/7.2 (1)
@@
Opus-4.6 Grok4 GPT-5.2 Gemini-3.1 Gemini-3 Grok4.1f DSR Kimi-K2t Hermes4 GPT-4o
T1 1.0/2.5 (5) 1.3/3.2 (4) 0.9/3.3 (2) 1.0/3.0 (3) 2.3/2.4 (3) 1.4/3.0 (5) 1.4/3.0 (5) 1.3/2.9 (4)
T2 1.2/2.7 (38) 1.5/2.9 (56) 1.4/2.8 (52) 1.5/2.9 (56) 1.3/2.4 (28) 1.6/3.0 (67) 1.6/2.9 (63) 1.7/3.0 (54) 1.8/3.0 (2) 2.9/3.2 (9)
T3 1.2/2.1 (36) 1.5/2.7 (42) 1.3/2.4 (34) 1.3/2.3 (38) 1.2/2.0 (22) 1.6/2.6 (49) 1.5/2.5 (43) 1.7/2.6 (34) 1.7/2.6 (24)
T4 0.9/1.7 (22) 1.6/3.0 (43) 1.4/2.7 (48) 1.5/2.3 (37) 1.3/2.4 (21) 2.0/3.1 (57) 1.7/2.9 (63) 1.9/3.3 (41) 5.0/5.6 (4)
T5 0.6/1.4 (21) 1.4/2.4 (27) 0.9/1.9 (27) 1.0/1.9 (31) 0.9/2.1 (21) 1.4/2.4 (28) 1.2/2.1 (33) 1.3/2.4 (24) 5.8/— (0)
@@
Opus-4.6 Grok4 GPT-5.2 Gemini-3.1 Gemini-3 Grok4.1f DSR Kimi-K2t Hermes4 GPT-4o
T1 1.0/1.8 (10) 1.2/2.4 (4) 1.4/2.3 (14) 1.2/2.3 (10) 1.0/2.0 (10) 1.8/2.0 (13) 1.8/2.5 (14) 2.7/2.3 (18) 4.0/4.0 (1) 5.5/9.2 (1)
T2 1.1/2.6 (14) 1.4/2.8 (11) 1.8/3.2 (14) 1.5/2.9 (16) 1.2/2.1 (14) 1.8/2.4 (16) 2.0/3.4 (16) 3.4/4.4 (18) 1.9/5.8 (1) 2.6/4.9 (7)
T3 1.4/2.4 (13) 1.3/2.4 (10) 1.7/3.1 (12) 1.4/2.2 (10) 1.2/1.8 (11) 1.8/2.7 (14) 1.7/2.7 (14) 3.2/4.3 (16) 2.0/4.4 (1) 2.8/4.3 (5)
T4 0.9/1.7 (13) 1.2/2.6 (11) 2.1/2.7 (17) 1.5/2.0 (15) 1.7/2.4 (14) 1.5/1.9 (8) 2.3/2.9 (24) 2.9/3.5 (19)
T5 1.1/1.0 (13) 1.6/1.3 (12) 2.4/1.3 (22) 1.7/1.2 (15) 1.8/1.3 (18) 2.0/1.6 (15) 2.5/1.5 (21) 2.5/1.5 (19) 4.9/4.2 (2) 6.9/— (0)
T6 0.2/0.1 (3) 0.2/0.1 (2) 0.2/0.7 (4) 0.2/0.1 (2) 0.2/0.3 (4) 0.5/1.3 (7) 0.3/0.8 (5) 0.2/0.0 (2) 2.9/3.2 (12)
T7 0.1/0.4 (13) 0.3/0.5 (12) 0.4/0.7 (16) 0.4/0.7 (16) 1.2/0.6 (12) 0.7/0.9 (19) 0.7/1.0 (19) 1.2/1.4 (21) 1.0/1.8 (1) 1.7/2.1 (15)
@@
(i) Hermes4 and GPT-4o achieve near-zero train validity for most theories in ABD-Full and ABD-Partial, reflecting their difficulty with multi-world abduction. (ii) In ABD-Skeptical, holdout gaps can be lower than train gaps (T5), because worst-case cost under universal completion varies across world draws. (iii) Opus-4.6 achieves the highest survivor counts and near-100% train validity across all scenarios, with competitive gap levels; in ABD-Skeptical it achieves the tightest gap distribution (median 0.20, only 9.4% above gap 3).
@@ G.2 Gap Distribution Summaries
Table G.4: Normalised train-gap distribution by scenario. N = predictions with all-valid on train. >>3, >>5 = tail percentages.
@@
Gemini-3 151 1.24 1.12 2.10 8.2 1.3% 0.7%
Grok4.1f 183 1.35 1.30 2.17 3.0 0.0% 0.0%
DSR 183 1.45 1.40 2.20 3.2 0.5% 0.0%
Kimi-K2t 111 1.48 1.50 2.20 2.9 0.0% 0.0%
Hermes4 6 2.85 2.25 5.80 5.8 16.7% 16.7%
GPT-4o 17 3.05 1.88 5.90 6.5 41.2% 35.3%
@@
Grok4 220 1.54 1.56 2.33 5.4 1.8% 0.5%
GPT-5.2 226 1.30 1.22 2.11 3.0 0.0% 0.0%
Gemini-3.1 236 1.41 1.22 2.13 7.0 2.5% 0.8%
Gemini-3 160 1.22 1.11 1.89 7.8 2.5% 0.6%
Grok4.1f 231 1.70 1.62 2.67 5.9 3.9% 0.4%
DSR 234 1.54 1.56 2.33 3.5 0.9% 0.0%
Kimi-K2t 176 1.69 1.62 2.62 4.0 3.4% 0.0%
Hermes4 2 1.78 1.78 2.33 2.3 0.0% 0.0%
GPT-4o 55 3.19 2.33 5.75 6.8 41.8% 23.6%
@@
DSR 151 1.73 1.33 4.40 6.8 21.9% 4.6%
Kimi-K2t 130 2.46 2.18 4.71 10.0 28.5% 6.2%
Hermes4 6 3.11 2.50 6.80 6.8 33.3% 16.7%
@@ Discussion.
In ABD-Full, the top six models show tightly concentrated gap distributions (P90 $\leq 2.3$, tail >3>3 at most 1.3%). ABD-Partial shows slightly heavier tails due to cost computation under existential completion. ABD-Skeptical exhibits the widest gap distributions: weaker models have over 20% of predictions with gap >3>3 (worst-case cost under universal completion inflates gaps), while Opus-4.6 achieves the tightest distribution (median 0.20, P90 2.83, only 9.4% above gap 3).

### v2

@@ 6.5 Complexity and Generalization
Grok4.1f 99 +1.4 64 +0.7 89 +0.1
DSR 96 +1.6 57 +0.7 75 +0.1
GPT-5.4 100 +1.9 77 +0.4 15 +0.6
Grok4 89 +1.5 69 +0.7 50 +0.8
Gemini-3 93 +1.3 61 +1.1 50 +0.7
Kimi-K2t 100 +1.5 73 +0.5 70 +1.8
@@
Grok4.1f 93 +1.2 33 -0.1 0
DSR 89 +1.1 70 +1.2
GPT-5.4 84 +1.0 32 +1.0 9 +0.8
Grok4 90 +1.1 17 +1.2 0
Gemini-3 82 +1.1 30 +0.4 27 +0.7
Kimi-K2t 95 +1.2 35 +0.9 0
@@
FULL 171 91 +1.47 171 35 +0.58
PARTIAL 193 82 +1.11 193 20 +0.59
SKEPTICAL 130 80 +0.04 130 31 +0.37
Overall 494 85 +0.95 494 28 +0.52
@@
Grok4.1f 7.0% 0.7 21.8 36
DSR 5.3% 0.7 19.6 29
GPT-5.4 56.8% 0.5 90.1 290
Grok4 11.7% 0.6 30.0 59
Gemini-3 11.7% 0.7 23.7 51
Kimi-K2t 3.5% 0.6 16.3 15
@@
Table 11 shows that some models beat the planted gold rule on a nontrivial fraction of instances, confirming that gold rules are generator anchors (Section 4.4). The clearest example is GPT-5.4, which dominates the beats-gold table but does so with very large formulas and weak holdout behavior. This diagnostic identifies instances where the planted rule may be suboptimal and feeds back into generator refinement.
@@ 6.6 Discussion
Models differ qualitatively in their performance profiles. Opus-4.6, Gemini-3.1, DSR, and Grok4.1f form a high-validity cluster; within that group, DSR and Grok4.1f are the strongest compact generalizers on holdout, with average valid-formula AST near 11, while Opus-4.6 is the train-validity leader. GPT-5.4 is the clearest low-gap outlier: it achieves the best train Gap and GapG, but does so with the largest valid formulas (AST 65.9 overall) and weak holdout survival (24.8% overall; 28.5% conditional on train-validity). Kimi-K2t shows nearly the opposite tradeoff, remaining compact and conditionally robust but with substantially larger gaps. Taken together, the results support three conclusions.
@@
Among the high-validity systems, cost parsimony remains challenging: train-set gap stays around one extra abnormality per world beyond the solver lower bound, and this gap roughly doubles on holdouts in ABD-Full/Partial.
@@
[…]
@@
[…]
@@ 7 Conclusion
We introduced ABD, a solver-checkable benchmark suite for default–exception abduction over finite first-order worlds. In ABD, models must synthesize a first-order abnormality definition $\alpha(x)$ such that $Ab(x)\leftrightarrow\alpha(x)$ repairs a fixed default theory across multiple worlds, while keeping exceptions sparse. We formalized three observation regimes—ABD-Full (closed-world), ABD-Partial (existential completion with best-case cost), and ABD-Skeptical (universal completion with worst-case cost)—and evaluated eleven LLMs under exact SMT-based verification.
Across tasks, frontier models often produce valid repairs on training worlds, but ABD remains unsaturated because validity, cost parsimony, and syntactic complexity do not saturate together. Among the high-validity models, the normalized gap to the solver lower bound remains roughly 1.01.01.61.6 extra abnormalities per world, and the more compact systems stay in the low teens of valid-formula AST, while GPT-5.4 reaches much lower train gaps only by using much larger formulas. The benchmark was calibrated to be moderately challenging for frontier models; harder instances can readily be generated in the future.
@@
[…]
@@
Software and Data
Acknowledgements
This work used AI assistants, including Anthropic Claude Code and OpenAI Codex, for software development, drafting, and editorial improvements to the manuscript and scripts. All substantive scientific decisions, experiment design and execution, result verification, and final writing were performed and validated by the authors. The authors remain fully responsible for the content of this submission.
@@ G.1 Full Per-Theory Results
Opus-4.6 GPT-5.2 Gemini-3.1 Grok4.1f DSR GPT-5.4 Grok4 Gemini-3 Kimi-K2t Hermes4 GPT-4o
T1 1.3/3.2 (4) 1.3/3.7 (3) 1.5/3.2 (4) 1.4/3.2 (5) 1.5/3.3 (5) 0.1/0.4 (1) 1.1/3.2 (3) 1.1/1.9 (3) 1.5/3.1 (4)
T2 1.4/3.2 (56) 1.1/3.2 (34) 1.4/3.2 (54) 1.4/2.9 (44) 1.6/3.6 (44) 0.3/1.3 (16) 0.8/1.5 (31) 1.2/2.8 (36) 1.3/3.5 (33) 2.0/5.3 (6)
T3 1.4/2.8 (37) 0.9/2.6 (17) 1.2/2.5 (39) 1.5/2.8 (39) 1.5/2.8 (34) 0.2/1.2 (7) 1.1/2.6 (27) 1.4/2.5 (19) 1.5/2.9 (26) 1.4/3.6 (4)
T4 1.3/2.3 (46) 0.9/2.4 (22) 1.4/2.1 (44) 1.3/2.3 (45) 1.5/2.6 (41) 0.0/0.4 (6) 0.9/1.4 (25) 1.4/2.2 (21) 1.7/2.6 (24) 5.3/6.4 (0)
T5 0.8/1.8 (27) 0.6/2.2 (17) 0.9/2.0 (27) 0.9/2.0 (23) 0.9/1.9 (27) 0.1/1.1 (4) 0.7/1.9 (21) 0.9/2.0 (24) 1.3/2.3 (24) 2.8/5.5 (5) 5.8/7.1 (1)
@@
Opus-4.6 GPT-5.2 Gemini-3.1 Grok4.1f DSR GPT-5.4 Grok4 Gemini-3 Kimi-K2t Hermes4 GPT-4o
T1 1.0/2.5 (5) 0.9/3.3 (2) 1.0/3.0 (3) 1.4/3.0 (5) 1.4/3.0 (5) 0.5/2.2 (2) 1.3/3.1 (4) 2.3/2.4 (3) 1.3/2.9 (4)
T2 1.2/2.6 (39) 1.4/2.8 (49) 1.5/2.8 (56) 1.6/3.0 (63) 1.6/2.9 (59) 0.7/2.6 (21) 1.6/2.9 (52) 1.3/2.3 (26) 1.7/3.0 (54) 1.8/3.0 (2) 2.9/3.2 (9)
T3 1.2/2.1 (36) 1.3/2.5 (32) 1.3/2.3 (38) 1.6/2.7 (46) 1.5/2.6 (40) 0.8/2.6 (11) 1.5/2.7 (39) 1.2/2.1 (20) 1.7/2.6 (34) 1.7/2.6 (24)
T4 0.9/1.7 (23) 1.4/2.6 (44) 1.5/2.3 (37) 2.0/3.1 (52) 1.7/2.9 (57) 0.6/1.8 (12) 1.7/2.9 (39) 1.3/2.2 (18) 1.9/3.2 (41) 5.0/5.8 (4)
T5 0.6/1.4 (21) 0.9/1.8 (25) 1.0/1.9 (31) 1.4/2.4 (26) 1.2/2.2 (30) 0.4/1.3 (12) 1.4/2.6 (26) 0.9/2.1 (19) 1.3/2.3 (25) 5.8/6.6 (0)
@@
Opus-4.6 GPT-5.2 Gemini-3.1 Grok4.1f DSR GPT-5.4 Grok4 Gemini-3 Kimi-K2t Hermes4 GPT-4o
T1 1.0/1.8 (10) 1.4/2.3 (14) 1.2/2.3 (10) 1.8/2.0 (13) 1.8/2.5 (14) 0.2/1.6 (4) 1.2/2.4 (4) 1.0/2.0 (10) 2.5/2.3 (18) 4.0/4.0 (1) 5.5/9.2 (1)
T2 1.1/2.6 (14) 1.8/3.2 (14) 1.5/2.9 (16) 1.8/2.4 (16) 1.8/3.4 (16) 0.6/0.7 (8) 1.4/2.8 (11) 1.2/2.1 (14) 3.3/4.4 (18) 1.9/5.8 (1) 2.6/4.9 (7)
T3 1.4/2.4 (13) 1.7/3.1 (12) 1.4/2.2 (10) 1.8/2.7 (14) 1.7/2.7 (14) 0.2/0.5 (6) 1.3/2.4 (10) 1.2/1.8 (11) 3.2/4.3 (16) 2.0/4.4 (1) 2.8/4.3 (5)
T4 0.9/1.7 (13) 2.1/2.7 (17) 1.5/2.0 (15) 1.5/1.9 (8) 2.3/2.9 (24) 0.6/1.0 (12) 1.2/2.6 (11) 1.7/2.6 (14) 2.8/3.3 (20)
T5 1.1/1.0 (13) 2.4/1.3 (22) 1.7/1.2 (15) 2.0/1.6 (15) 2.5/1.5 (21) 0.8/0.9 (11) 1.6/1.3 (12) 1.7/1.3 (18) 2.5/1.5 (19) 4.9/4.2 (2) 6.9/— (0)
T6 0.2/0.1 (3) 0.2/0.7 (4) 0.2/0.1 (2) 0.4/1.3 (7) 0.3/0.8 (5) 0.0/0.3 (3) 0.2/0.1 (2) 0.2/0.2 (4) 0.2/0.5 (2) 2.9/3.2 (12)
T7 0.1/0.4 (13) 0.3/0.7 (16) 0.3/0.7 (16) 0.7/0.9 (19) 0.7/1.0 (19) 0.0/0.1 (9) 0.3/0.5 (12) 1.2/0.6 (12) 1.3/1.4 (21) 1.0/1.8 (1) 1.7/2.1 (15)
@@
(i) Hermes4 and GPT-4o achieve near-zero train validity for most theories in ABD-Full and ABD-Partial, reflecting their difficulty with multi-world abduction. (ii) In ABD-Skeptical, holdout gaps can be lower than train gaps (T5), because worst-case cost under universal completion varies across world draws. (iii) Opus-4.6 combines the highest survivor counts with near-perfect train validity across all scenarios, while GPT-5.4 is the clearest cost outlier: many per-theory train gaps are near zero, but survivor counts are much smaller.
@@ G.2 Gap Distribution Summaries
Table G.4: Normalised train-gap distribution by scenario. N = predictions with all-valid on train. >>3 and >>5 are tail percentages.
@@
Grok4.1f 171 1.32 1.25 2.12 3.0 0.0% 0.0%
DSR 172 1.42 1.39 2.16 3.2 0.6% 0.0%
GPT-5.4 149 0.13 0.00 0.38 2.4 0.0% 0.0%
Grok4 156 0.90 0.89 1.62 2.4 0.0% 0.0%
Gemini-3 138 1.23 1.12 2.00 8.2 1.4% 0.7%
Kimi-K2t 117 1.45 1.40 2.18 2.9 0.0% 0.0%
Hermes4 5 2.84 2.20 4.40 5.8 20.0% 20.0%
GPT-4o 17 3.05 1.88 5.84 6.5 41.2% 35.3%
@@
GPT-5.2 208 1.30 1.22 2.11 3.0 0.0% 0.0%
Gemini-3.1 236 1.40 1.22 2.13 7.0 2.5% 0.8%
Grok4.1f 213 1.69 1.62 2.61 5.9 3.3% 0.5%
DSR 216 1.54 1.56 2.33 3.5 0.9% 0.0%
GPT-5.4 210 0.65 0.50 1.62 3.0 0.0% 0.0%
Grok4 205 1.55 1.56 2.33 5.4 2.0% 0.5%
Gemini-3 147 1.22 1.11 1.93 7.8 2.7% 0.7%
Kimi-K2t 178 1.68 1.62 2.58 4.0 3.4% 0.0%
Hermes4 2 1.78 1.78 2.22 2.3 0.0% 0.0%
GPT-4o 55 3.18 2.33 5.72 6.8 41.8% 23.6%
@@
DSR 155 1.69 1.29 4.37 6.8 21.3% 4.5%
GPT-5.4 152 0.37 0.00 0.98 5.2 3.9% 0.7%
Grok4 144 1.08 0.33 3.12 6.7 10.4% 2.1%
Gemini-3 150 1.25 0.43 3.60 6.7 15.3% 4.7%
Kimi-K2t 134 2.41 2.08 4.64 10.0 27.6% 6.0%
Hermes4 6 3.11 2.50 5.40 6.8 33.3% 16.7%
@@ Discussion.
In ABD-Full, the high-validity models still show fairly tight train-gap distributions, with P90 usually around 2.0–2.2 and almost no mass above gap 3. ABD-Partial is slightly heavier-tailed due to cost computation under existential completion. GPT-5.4 is the dominant low-gap outlier in every scenario (mean 0.13 in ABD-Full, 0.65 in ABD-Partial, 0.37 in ABD-Skeptical), but this concentration is a train-set parsimony effect rather than a sign of strong holdout robustness. Among the remaining high-validity models, ABD-Skeptical exhibits the widest tails: Opus-4.6 still has the tightest skeptical distribution (median 0.18, only 9.4% above gap 3), while GPT-5.2, DSR, and Kimi-K2t all exceed 20% above gap 3.